OpenStudy (anonymous):

Hi. I wanted to know how does one know when to use the substitution rule for integrals and when not too?

OpenStudy (saifoo.khan):

@satellite73

OpenStudy (lgbasallote):

take the derivative of u....if that du is present in the integrand you can use sub

OpenStudy (anonymous):

don't use them on saturday night. go out and party

OpenStudy (lgbasallote):

wait...we're talking about u-sub right? not algebraic sub?

OpenStudy (lgbasallote):

lol sat

OpenStudy (anonymous):

@lgbasallote yep I am. I wanted to know when you look at an integral problem, how do you know to use either u-substitution or not.

OpenStudy (lgbasallote):

go out and integrate?

OpenStudy (lgbasallote):

@malexander it involves mastery of derivatives...if you know many derivativves you can see them immediately...if not..you can try u-ing then derive it...

OpenStudy (anonymous):

a function multiplied by something that is either the derivative, or a constant multiple of the derivative

OpenStudy (anonymous):

\[\int\sqrt{x^3+3}x^2dx\] for example, because the derivative of \(x^3+3\) is \(3x^2\) and what i wrote above was wrong

OpenStudy (anonymous):

should have been \[\int f(g(x))g'(x)dx\]

OpenStudy (lgbasallote):

^integral of the chain rule

OpenStudy (lgbasallote):

just noticed lo

OpenStudy (lgbasallote):

lol*

OpenStudy (anonymous):

Don't use it on Sundays either, yolo.

OpenStudy (anonymous):

Your confusing me @satellite73. Um. Lets say you have these four integrals (question 2). In this question, which one would you need u-substitution and why?

OpenStudy (lgbasallote):

2c

OpenStudy (lgbasallote):

2b

OpenStudy (anonymous):

Why those @lgbasallote ?

OpenStudy (lgbasallote):

\[\large \int \frac{3x^3}{\sqrt{x^4 + 1}} = 3\int \frac{x^3}{\sqrt{x^4 + 1}}\] if i let u = x^4 + 1 du = 4x^3 the constant 4 is not important...but x^3 is...and it is present in the integrand so you can use u-sub

OpenStudy (lgbasallote):

2d is u-subbable too

OpenStudy (anonymous):

Oh okay. That makes a bit more sense @lgbasallote . However, I thought we were not able to integrate fractions? (Im a cal 1 student, maybe inter grating fractions is in cal 2)

OpenStudy (lgbasallote):

integrating fractions....for example 1/u^2 1/u^2 = u^(-2) according to basic algebra according to basic integration...you can use power rule on that..i assume you're familiar with that note: NEVER use power rule on 1/u <---the integral of that is automatically ln u

OpenStudy (lgbasallote):

also note that i am using u...this is because u is a function of x meaning it is an expression..not limited to only a variable

OpenStudy (anonymous):

right, i remember that @lgbasallote about 1/u^2..and also thats fine, we use "u" in class

OpenStudy (anonymous):

@lgbasallote ln is in calc 2, so i hope i get to do that next semester hehe...i think i just need to do more problems with integration then...

OpenStudy (lgbasallote):

i see...so you're just learning u-sub...try integrating the ones i said..2b and 2d

OpenStudy (anonymous):

will do!

OpenStudy (anonymous):

thanks again @lgbasallote ....i have another question...do u mind me asking it here, or posting another question box?

OpenStudy (lgbasallote):

new post...im too lagged for a long thread lol

OpenStudy (anonymous):

LOL! ok sounds good